# Signed Networks: theory, methods, and applications

### v1

@@ 2.2.4 Signed centralities
Centrality is a fundamental concept in network analysis, widely used to identify the most structurally important nodes in a network. Classical measures such as degree, closeness, betweenness, and eigenvector centrality have been extensively studied in unsigned networks. Central nodes can be defined as those with many connections (degree centrality), those who are at a short distance from others (closeness centrality), those who lie on many shortest paths (betweenness centrality), or, in a circular way, as those with connections with many other central nodes (eigenvector centrality). These measures can often be interpreted in terms of network flow processes [borgattiCentralityNetworkFlow2005]. In signed networks, the notion of centrality becomes more nuanced, as flow-based interpretations become less applicable. Depending on the context, centrality in signed networks can represent many concepts, such as influence, power, trustworthiness, control, cohesiveness, or balance.
@@
The simplest approach is to extend classical centrality measures by treating positive and negative links equivalently. This perspective views centrality as a measure of structural importance without distinguishing between different types of relationships. A similar approach, used for instance in P2P networks [kamvarEigenTrustAlgorithmReputation], is to ignore negative links entirely and compute classical centrality measures on the positive subgraph. This assumption fails to distinguish between a negative link and the absence of a link, discarding structural information. Another approach computes centralities separately for positive and negative subgraphs and subtracts them [shahriariRankingNodesSigned2014]. While this accounts for negative links, it ignores their interactions with positive ones, leading to outputs that are often difficult to interpret.
@@
Thus, these straightforward adaptations of unsigned centrality measures often fail to provide interpretable results. A simpler way forward is to start by asking what kind of importance we want to capture, rather than adapting existing formulas. Once this notion is clear, suitable mathematical definitions can follow. For instance, in politically charged networks, centrality can represent power or control, and the mathematical formulation changes depending on which notion one aims to capture [smithPowerPoliticallyCharged2014]. If we interpret centrality as power, positive edges increase it by reinforcing influence and access to allies, while negative edges reduce it by introducing threats and opposition. If, instead, we view centrality as control, the effect reverses: a node may gain centrality when its allies face many negative edges, as they rely more on it for stability and mediation [smithPowerPoliticallyCharged2014].
@@
Other perspectives of centrality in signed networks include measures quantifying trust or cohesiveness. Trust is particularly relevant in social networks, recommendation systems, and online marketplaces, where an individual’s credibility is key [guhaPropagationTrustDistrust2004, tangSurveySignedNetwork2016]. Trust-based centrality measures emphasise the asymmetry between trust and distrust, often weighing negative endorsements more heavily [guhaPropagationTrustDistrust2004]. In some other applications, being central means being embedded in a cohesive cluster of mutual support [bonacichCalculatingStatusNegative2004, gromovSocialBalancebasedCentrality2025]. In this framework, negative links can increase centrality by reinforcing cohesion within a group through shared adversaries. Centrality then reflects how well an individual contributes to the stability of a coalition rather than mere connectivity. Finally, centrality in signed networks can also be linked to balance theory (see Chapter 3). Some classical measures, like subgraph centrality, naturally function as local balance indicators when applied to signed networks [diaz-diazMathematicalModelingLocal2024].
@@
Since shortest paths are not well-defined in signed networks, closeness centrality relies on alternative distance measures. One approach replaces the shortest-path distance with a generalised distance. The generalised closeness centrality is then [estradaCommunicabilityCosineDistance2024]
@@
Katz centrality extends degree centrality by considering both direct and indirect connections, penalising walks of length $\ell$ by an attenuation factor $\alpha^{\ell}$ ($\alpha<1$). In unsigned networks, it is defined as [newmanNetworks2018] [^5] The Katz centrality is only well-defined when $\alpha$ is smaller than the inverse of the spectral radius of $A$.
@@
where $C_{K}(i)$ counts all walks of all lengths $\ell$ starting from node $i$, with more importance given to the shortest ones. In signed networks, Katz centrality was first introduced in [chiangExploitingLongerCycles2011] and later formalised in [gomez-ambrosiMeasuringNetEffects2025]. The same formula applies, but its interpretation differs: in unsigned networks, all walks contribute positively to a node’s influence, whereas in signed ones, walks with an odd number of negative edges contribute negatively. The resulting value thus reflects the net level of power of a node. Unlike in the unsigned case, Katz centrality can take negative values, indicating that the node has, on average, stronger antagonistic relations. Variants with different weightings on positive and negative edges have also been explored [everettNetworksContainingNegative2014].
@@
The core idea of eigenvector centrality is that a node is considered important if it is connected to other important nodes [bonacichUniquePropertiesEigenvector2007]. In the signed case, the most natural generalisation assumes that a node’s importance increases when it is connected to important nodes through positive edges, and decreases when connected via negative edges to important nodes. Mathematically, this is captured by the following self-referential equation:
@@
In signed networks, the interpretation of eigenvector centrality depends critically on whether the network is structurally balanced (see Section 3.4). In fact, in a balanced network, the adjacency matrix is guaranteed to have a dominant eigenvector associated with the largest eigenvalue, which will be non-degenerate [bonacichCalculatingStatusNegative2004]. Unlike in the unsigned case, the components of this eigenvector can take both positive and negative values, and their signs align with the two balanced factions of the network. Eigenvector centrality can thus be interpreted as a measure of a node’s adherence to its faction, with the magnitude reflecting its structural importance and the sign indicating its faction.
When the network is unbalanced, the largest eigenvalue can become degenerate, and the dominant eigenvector is no longer uniquely defined. Everett et al. [everettNetworksContainingNegative2014] illustrate this issue with two signed networks whose eigenvector centralities are ill-defined. Even so, the measure often gives useful results when the network is close to being balanced [diazdiazthesis] (see Section 3.5).
@@
Originally designed by Google for ranking webpages [pagePagerankCitationRanking1999], PageRank centrality is akin to a directed version of eigenvector centrality that incorporates a damping factor $\alpha$ to ensure that every node receives a minimal level of importance, even if it has no incoming edges. While eigenvector centrality can be interpreted as the stationary distribution of a random walk on the graph, PageRank centrality is the stationary distribution of a random walk with teleportations happening at rate $\alpha$. In unsigned networks, PageRank satisfies the recursive equation:
@@
where $k_{j}^{out}$ is the out-degree of node $j$. This formulation captures the intuition that a node’s importance is determined by the importance of the nodes linking to it, weighted by their out-degree. In signed networks, the most natural generalisation replaces $A$ with the signed adjacency matrix while still using the underlying out-degree [kunegisSlashdotZooMining2009]. This minimal modification keeps the core properties of PageRank while taking into account that negative links penalise centrality. However, note that several properties get lost when including negative edges. For example, some PageRank values may become negative, so they can no longer be interpreted as a probability distribution.
@@
Beyond adaptations of classical measures, several alternative approaches to centrality in signed networks have been proposed. Approaches considering signed links as indicators of trust or distrust include PageTrust [kerchovePageTrustAlgorithmHow2008], inspired by PageRank, and exponential ranking [traagExponentialRankingTaking2010]. Bias and prestige centrality assume that the trustworthiness and reputation of a node should be assessed separately [mishraFindingBiasPrestige2011], much like the HITS algorithm [kleinbergWebGraphMeasurements1999]. Finally, Liu et al. proposed a framework reminiscent of signed PageRank, but distinguishing between a node’s total centrality (i.e. its overall influence on the network), and its net centrality (which reveals whether that influence is mainly positive or negative) [liuSimpleApproachQuantifying2020].
@@ 2.2.5 Signed motifs
A network motif is a recurring subgraph pattern that appears significantly more frequently in a given network than in randomised counterparts [miloNetworkMotifsSimple2002]. These motifs are fundamental building blocks of complex networks and provide insights into their structural and functional properties. The detection of motifs typically involves searching for all possible small subgraphs, such as triads (three-node subgraphs) or tetrads, and comparing their frequency to a null model to ensure statistical significance. One of the simplest ways to quantify this significance is through the Z-score:
@@
Despite these challenges, studies have identified and classified significant motif patterns in signed networks. Two complementary frameworks are commonly used to interpret these motifs: structural balance and status theory. In the balance framework, signed motifs are divided into balanced and unbalanced types (see Chapter 3). The former contain only cycles with an even number of negative edges, whereas the latter include cycles with an odd number of negative edges. Analyses of empirical signed networks often report that nearly all balanced motifs involving three or four nodes are overrepresented, while nearly all unbalanced motifs are underrepresented [haoProperNetworkRandomization2023]. However, one must keep in mind that these statistical patterns depend strongly on the type of data and the choice of null model [galloTestingStructuralBalance2024], and the evidence across the literature remains mixed.
Moreover, status theory [leskovecSignedNetworksSocial2010] provides a hierarchical interpretation of signed motifs. It assumes that each node occupies a latent status level, and that the sign and direction of edges reflect relative positions within this hierarchy. A positive directed edge indicates that the source node perceives the target node as having a higher status, whereas a negative edge implies the opposite. Status theory explains asymmetries in social networks, such as the tendency of lower-status individuals to form positive links with higher-status individuals [leskovecSignedNetworksSocial2010].
@@ 2.3 Signed Laplacian
A key question in extending the Laplacian to signed networks is how to define the degree of each node. One option is to employ the underlying degree, while another is to use the net degree. The first choice results in a matrix commonly known as the opposing Laplacian [houLaplacianEigenvaluesSigned2003, houBoundsLeastLaplacian2005], while the latter leads to the repelling Laplacian [shiDynamicsSignedNetworks2019].
@@
This matrix has the advantage of being always positive semidefinite, meaning all its eigenvalues are non-negative [kunegisSpectralAnalysisSigned2010]. This ensures that any associated dynamics are stable (see Section 5.1). Moreover, it has a zero eigenvalue if and only if the graph is balanced. In such cases, the signs of the components of the eigenvector corresponding to the zero eigenvalue can be used to determine the balanced faction structure of the network [^6] A faction is a set of nodes whose mutual connections are mostly positive and whose connections to nodes outside the set are mostly negative. A balanced faction is an ideal case where every link within the set is positive and every link to other sets is negative. See Section 3.4 for more details. [kunegisSpectralAnalysisSigned2010]. This property makes the opposing Laplacian particularly useful for uncovering faction structures and community partitions in signed networks  [cucuringuSPONGEGeneralizedEigenproblem2019] (see Section 4.4) [^7] It is instructive to compare the spectrum of $L_{o}$ with that of $L_{u}$. The smallest eigenvalue of $L_{o}$ plays the same role as the second smallest eigenvalue of $L_{u}$: both vanish when the network is balanced or disconnected, respectively, and their associated eigenvectors uncover balanced factions or connected components. When these eigenvalues are small but nonzero, the network is nearly balanced or nearly disconnected, and the corresponding eigenvector reveals faction or community structures, respectively..
@@
This slight alteration shifts the focus from structural balance to the dynamical behaviour of the network [panLaplacianDynamicsSigned2016]. Like the unsigned Laplacian, the repelling Laplacian always has a zero eigenvalue, with its associated eigenvector being the constant vector. Unlike the opposing Laplacian, $L_{r}$ is not necessarily positive semidefinite. This means that the repelling Laplacian is better suited for capturing instabilities and repulsive interactions that may lead to divergent behaviour in the network.
Both of these variants find a wide range of applications. The opposing Laplacian is instrumental in tasks such as spectral embedding for node clustering, random walks that explore the network structure, and the analysis of generalised distances like resistance distance [tianSpreadingStructuralBalance2024, kunegisApplicationsStructuralBalance2014, altafiniConsensusProblemsNetworks2013]. On the other hand, the repelling Laplacian has been employed in applications that require an understanding of dynamical stability, such as the analysis of electrical and mechanical networks where interactions can be repulsive, as well as consensus and synchronisation models where the stability of the evolving state is a central concern [shiDynamicsSignedNetworks2019, fontanPseudoinversesSignedLaplacian2023].
@@
[ caption = Comparison between the unsigned, opposing and repelling Laplacians. $G$ is assumed to be undirected and connected., label = tab:laplacian, ] colspec = X X X X, width = cell11 = cmd=, cell12 = cmd=, cell13 = cmd=, cell14 = cmd=, Unsigned Laplacian, $L_{u}$ Opposing Laplacian, $L_{o}$ Repelling Laplacian, $L_{r}$
@@
Applications Node embeddings [vonLuxburg2007a], random walks [masuda2017random], linear consensus [degrootReachingConsensus1974], synchronization [arenasSynchronizationComplexNetworks2008], generalized distances (resistance distance) [devriendtEffectiveResistanceMore2022, chandraElectricalResistanceGraph1989], analysis of electrical and spring networks [chandraElectricalResistanceGraph1989, lovaszRandomWalksGraphs1993], kernel functions [foussRandomWalkComputationSimilarities2007]. Node embeddings [kunegisSpectralAnalysisSigned2010, cucuringuSPONGEGeneralizedEigenproblem2019], random walks [tianSpreadingStructuralBalance2024], linear consensus [altafiniConsensusProblemsNetworks2013], generalized distances (resistance distance) [kunegisApplicationsStructuralBalance2014], kernel functions [kunegisSpectralAnalysisSigned2010]. Node embedding (SHEEP) [babulSHEEPSignedHamiltonian2024], analysis of electrical and spring networks [fontanPseudoinversesSignedLaplacian2023], linear consensus, synchronization [panLaplacianDynamicsSigned2016, shiDynamicsSignedNetworks2019].
@@ 3 Cognitive and structural balance
In the previous chapter, we treated signed networks as a mathematical abstraction independent of the specific system they represent. This approach has the advantage of yielding general results applicable to a wide variety of systems, but it also forgets about the unique characteristics of those individual systems. Furthermore, it overlooks the historical origins of the field in the study of social systems, which influence the terminology, concepts and methods we use nowadays. This observation becomes particularly relevant when discussing the concept of balance. Balance is a central concept in signed networks that can be traced back to the origins of the field [heiderAttitudesCognitiveOrganization1946, cartwrightAmbivalenceIndifferenceGeneralizations1970]. In a nutshell, it refers to the way connections in a system are arranged to produce a sense of consistency or stability for the individuals that form that system. As an example, a configuration where the friend of my friend is my friend and the enemy of my enemy is my friend would be considered balanced.
The signed network literature contains two closely related yet distinct notions often referred to as “balance”: cognitive balance, introduced by Fritz Heider [heiderAttitudesCognitiveOrganization1946], and structural balance, introduced by Cartwright and Harary [hararyNotionBalanceSigned1953, cartwrightStructuralBalanceGeneralization1956]. Both frameworks aim to capture the idea that certain configurations of positive and negative relations are more stable or consistent than others; nevertheless, they diverge significantly in scope and interpretation. Heiderian balance, rooted in cognitive psychology, focuses on how individuals perceive and respond to patterns of relations in their immediate social environment. It emphasises subjective (local) experience and psychological tension as drivers of network change. In contrast, Hararian balance abstracts away individual perception to define balance as a global structural property of a signed network. This formalisation enables the application of graph-theoretical tools and extends the notion of balance beyond social contexts to any domain involving antagonistic interactions. This chapter compares these two perspectives, clarifying common misconceptions and identifying contexts where each framework is most applicable. It then surveys several mathematical properties of balanced networks. Finally, it introduces a range of balance indices to quantify Hararian balance in empirical data.

### v2

@@ 2.2.4 Signed centralities
Centrality is a fundamental concept in network analysis, widely used to identify the most structurally important nodes in a network. Classical measures such as degree, closeness, betweenness, and eigenvector centrality have been extensively studied in unsigned networks. Central nodes can be defined as those with many connections (degree centrality), those who are at a short distance from others (closeness centrality), those who lie on many shortest paths (betweenness centrality), or, in a circular way, as those with connections with many other central nodes (eigenvector centrality). These measures can often be interpreted in terms of network flow processes [47]. In signed networks, the notion of centrality becomes more nuanced, as flow-based interpretations become less applicable. Depending on the context, centrality in signed networks can represent many concepts, such as influence, power, trustworthiness, control, cohesiveness, or balance.
@@
The simplest approach is to extend classical centrality measures by treating positive and negative links equivalently. This perspective views centrality as a measure of structural importance without distinguishing between different types of relationships. A similar approach, used for instance in P2P networks [205], is to ignore negative links entirely and compute classical centrality measures on the positive subgraph. This assumption fails to distinguish between a negative link and the absence of a link, discarding structural information. Another approach computes centralities separately for positive and negative subgraphs and subtracts them [349]. While this accounts for negative links, it ignores their interactions with positive ones, leading to outputs that are often difficult to interpret.
@@
Thus, these straightforward adaptations of unsigned centrality measures often fail to provide interpretable results. A simpler way forward is to start by asking what kind of importance we want to capture, rather than adapting existing formulas. Once this notion is clear, suitable mathematical definitions can follow. For instance, in politically charged networks, centrality can represent power or control, and the mathematical formulation changes depending on which notion one aims to capture [357]. If we interpret centrality as power, positive edges increase it by reinforcing influence and access to allies, while negative edges reduce it by introducing threats and opposition. If, instead, we view centrality as control, the effect reverses: a node may gain centrality when its allies face many negative edges, as they rely more on it for stability and mediation [357].
@@
Other perspectives of centrality in signed networks include measures quantifying trust or cohesiveness. Trust is particularly relevant in social networks, recommendation systems, and online marketplaces, where an individual’s credibility is key [157, 379]. Trust-based centrality measures emphasise the asymmetry between trust and distrust, often weighing negative endorsements more heavily [157]. In some other applications, being central means being embedded in a cohesive cluster of mutual support [44, 155]. In this framework, negative links can increase centrality by reinforcing cohesion within a group through shared adversaries. Centrality then reflects how well an individual contributes to the stability of a coalition rather than mere connectivity. Finally, centrality in signed networks can also be linked to balance theory (see Chapter 3). Some classical measures, like subgraph centrality, naturally function as local balance indicators when applied to signed networks [90].
@@
Since shortest paths are not well-defined in signed networks, closeness centrality relies on alternative distance measures. One approach replaces the shortest-path distance with a generalised distance. The generalised closeness centrality is then [110]
@@
Katz centrality extends degree centrality by considering both direct and indirect connections, penalising walks of length $\ell$ by an attenuation factor $\alpha^{\ell}$ ($\alpha<1$). In unsigned networks, it is defined as [296] [^5] The Katz centrality is only well-defined when $\alpha$ is smaller than the inverse of the spectral radius of $A$.
@@
where $C_{K}(i)$ counts all walks of all lengths $\ell$ starting from node $i$, with more importance given to the shortest ones. In signed networks, Katz centrality was first introduced in [68] and later formalised in [149]. The same formula applies, but its interpretation differs: in unsigned networks, all walks contribute positively to a node’s influence, whereas in signed ones, walks with an odd number of negative edges contribute negatively. The resulting value thus reflects the net level of power of a node. Unlike in the unsigned case, Katz centrality can take negative values, indicating that the node has, on average, stronger antagonistic relations. Variants with different weightings on positive and negative edges have also been explored [111].
@@
The core idea of eigenvector centrality is that a node is considered important if it is connected to other important nodes [45]. In the signed case, the most natural generalisation assumes that a node’s importance increases when it is connected to important nodes through positive edges, and decreases when connected via negative edges to important nodes. Mathematically, this is captured by the following self-referential equation:
@@
In signed networks, the interpretation of eigenvector centrality depends critically on whether the network is structurally balanced (see Section 3.4). In fact, in a balanced network, the adjacency matrix is guaranteed to have a dominant eigenvector associated with the largest eigenvalue, which will be non-degenerate [44]. Unlike in the unsigned case, the components of this eigenvector can take both positive and negative values, and their signs align with the two balanced factions of the network. Eigenvector centrality can thus be interpreted as a measure of a node’s adherence to its faction, with the magnitude reflecting its structural importance and the sign indicating its faction.
When the network is unbalanced, the largest eigenvalue can become degenerate, and the dominant eigenvector is no longer uniquely defined. Everett et al. [111] illustrate this issue with two signed networks whose eigenvector centralities are ill-defined. Even so, the measure often gives useful results when the network is close to being balanced [93] (see Section 3.5).
@@
Originally designed by Google for ranking webpages [300], PageRank centrality is akin to a directed version of eigenvector centrality that incorporates a damping factor $\alpha$ to ensure that every node receives a minimal level of importance, even if it has no incoming edges. While eigenvector centrality can be interpreted as the stationary distribution of a random walk on the graph, PageRank centrality is the stationary distribution of a random walk with teleportations happening at rate $\alpha$. In unsigned networks, PageRank satisfies the recursive equation:
@@
where $k_{j}^{out}$ is the out-degree of node $j$. This formulation captures the intuition that a node’s importance is determined by the importance of the nodes linking to it, weighted by their out-degree. In signed networks, the most natural generalisation replaces $A$ with the signed adjacency matrix while still using the underlying out-degree [223]. This minimal modification keeps the core properties of PageRank while taking into account that negative links penalise centrality. However, note that several properties get lost when including negative edges. For example, some PageRank values may become negative, so they can no longer be interpreted as a probability distribution.
@@
Beyond adaptations of classical measures, several alternative approaches to centrality in signed networks have been proposed. Approaches considering signed links as indicators of trust or distrust include PageTrust [114], inspired by PageRank, and exponential ranking [388]. Bias and prestige centrality assume that the trustworthiness and reputation of a node should be assessed separately [275], much like the HITS algorithm [215]. Finally, Liu et al. proposed a framework reminiscent of signed PageRank, but distinguishing between a node’s total centrality (i.e., its overall influence on the network), and its net centrality (which reveals whether that influence is mainly positive or negative) [248].
@@ 2.2.5 Signed motifs
A network motif is a recurring subgraph pattern that appears significantly more frequently in a given network than in randomised counterparts [273]. These motifs are fundamental building blocks of complex networks and provide insights into their structural and functional properties. The detection of motifs typically involves searching for all possible small subgraphs, such as triads (three-node subgraphs) or tetrads, and comparing their frequency to a null model to ensure statistical significance. One of the simplest ways to quantify this significance is through the Z-score:
@@
Despite these challenges, studies have identified and classified significant motif patterns in signed networks. Two complementary frameworks are commonly used to interpret these motifs: structural balance and status theory. In the balance framework, signed motifs are divided into balanced and unbalanced types (see Chapter 3). The former contain only cycles with an even number of negative edges, whereas the latter include cycles with an odd number of negative edges. Analyses of empirical signed networks often report that nearly all balanced motifs involving three or four nodes are overrepresented, while nearly all unbalanced motifs are underrepresented [167]. However, one must keep in mind that these statistical patterns depend strongly on the type of data and the choice of null model [138], and the evidence across the literature remains mixed.
Moreover, status theory [232] provides a hierarchical interpretation of signed motifs. It assumes that each node occupies a latent status level, and that the sign and direction of edges reflect relative positions within this hierarchy. A positive directed edge indicates that the source node perceives the target node as having a higher status, whereas a negative edge implies the opposite. Status theory explains asymmetries in social networks, such as the tendency of lower-status individuals to form positive links with higher-status individuals [232].
@@ 2.3 Signed Laplacian
A key question in extending the Laplacian to signed networks is how to define the degree of each node. One option is to employ the underlying degree, while another is to use the net degree. The first choice results in a matrix commonly known as the opposing Laplacian [187, 186], while the latter leads to the repelling Laplacian [353].
@@
This matrix has the advantage of being always positive semidefinite, meaning all its eigenvalues are non-negative [224]. This ensures that any associated dynamics are stable (see Section 5.1). Moreover, it has a zero eigenvalue if and only if the graph is balanced. In such cases, the signs of the components of the eigenvector corresponding to the zero eigenvalue can be used to determine the balanced faction structure of the network [^6] A faction is a set of nodes whose mutual connections are mostly positive and whose connections to nodes outside the set are mostly negative. A balanced faction is an ideal case where every link within the set is positive and every link to other sets is negative. See Section 3.4 for more details. [224]. This property makes the opposing Laplacian particularly useful for uncovering faction structures and community partitions in signed networks  [77] (see Section 4.4) [^7] It is instructive to compare the spectrum of $L_{o}$ with that of $L_{u}$. The smallest eigenvalue of $L_{o}$ plays the same role as the second smallest eigenvalue of $L_{u}$: both vanish when the network is balanced or disconnected, respectively, and their associated eigenvectors uncover balanced factions or connected components. When these eigenvalues are small but nonzero, the network is nearly balanced or nearly disconnected, and the corresponding eigenvector reveals faction or community structures, respectively..
@@
This slight alteration shifts the focus from structural balance to the dynamical behaviour of the network [302]. Like the unsigned Laplacian, the repelling Laplacian always has a zero eigenvalue, with its associated eigenvector being the constant vector. Unlike the opposing Laplacian, $L_{r}$ is not necessarily positive semidefinite. This means that the repelling Laplacian is better suited for capturing instabilities and repulsive interactions that may lead to divergent behaviour in the network.
Both of these variants find a wide range of applications. The opposing Laplacian is instrumental in tasks such as spectral embedding for node clustering, random walks that explore the network structure, and the analysis of generalised distances like resistance distance [383, 225, 9]. On the other hand, the repelling Laplacian has been employed in applications that require an understanding of dynamical stability, such as the analysis of electrical and mechanical networks where interactions can be repulsive, as well as consensus and synchronisation models where the stability of the evolving state is a central concern [353, 122].
@@
[ caption = Comparison between the unsigned, opposing, and repelling Laplacians. $G$ is assumed to be undirected and connected., label = tab:laplacian, ] colspec = X X X X, width = cell11 = cmd=, cell12 = cmd=, cell13 = cmd=, cell14 = cmd=, Unsigned Laplacian, $L_{u}$ Opposing Laplacian, $L_{o}$ Repelling Laplacian, $L_{r}$
@@
Applications Node embeddings [115], random walks [264], linear consensus [81], synchronization [23], generalized distances (resistance distance) [87, 61], analysis of electrical and spring networks [61, 250], kernel functions [126]. Node embeddings [224, 77], random walks [383], linear consensus [9], generalized distances (resistance distance) [225], kernel functions [224]. Node embedding (SHEEP) [28], analysis of electrical and spring networks [122], linear consensus, synchronization [302, 353].
@@ 3 Cognitive and structural balance
In the previous chapter, we treated signed networks as a mathematical abstraction independent of the specific system they represent. This approach has the advantage of yielding general results applicable to a wide variety of systems, but it also forgets about the unique characteristics of those individual systems. Furthermore, it overlooks the historical origins of the field in the study of social systems, which influence the terminology, concepts, and methods we use nowadays. This observation becomes particularly relevant when discussing the concept of balance. Balance is a central concept in signed networks that can be traced back to the origins of the field [177, 56]. In a nutshell, it refers to the way connections in a system are arranged to produce a sense of consistency or stability for the individuals that form that system. As an example, a configuration where the friend of my friend is my friend and the enemy of my enemy is my friend would be considered balanced.
The signed network literature contains two closely related yet distinct notions often referred to as “balance”: cognitive balance, introduced by Fritz Heider [177], and structural balance, introduced by Cartwright and Harary [168, 55]. Both frameworks aim to capture the idea that certain configurations of positive and negative relations are more stable or consistent than others; nevertheless, they diverge significantly in scope and interpretation. Heiderian balance, rooted in cognitive psychology, focuses on how individuals perceive and respond to patterns of relations in their immediate social environment. It emphasises subjective (local) experience and psychological tension as drivers of network change. In contrast, Hararian balance abstracts away individual perception to define balance as a global structural property of a signed network. This formalisation enables the application of graph-theoretical tools and extends the notion of balance beyond social contexts to any domain involving antagonistic interactions. This chapter compares these two perspectives, clarifying common misconceptions and identifying contexts where each framework is most applicable. It then surveys several mathematical properties of balanced networks. Finally, it introduces a range of balance indices to quantify Hararian balance in empirical data.